What is the pH of a buffer solution that is 0.400 M in HF and 0.790 M in NaF? Ka of HF = 7.2 ✕ 10-4.
Answer #1

PH   = PKa + log[NaF]/[HF]

Ka    = 7.2*10-4

PKa   = -logKa

        = -log7.2*10-4

        = 3.1426

PH    = 3.1426 + log0.79/0.4

       = 3.1426 + 0.2955   = 3.4381 >>>>> answer